# Context Parametrization with Compositional Adapters

### v2

@@ Abstract
Large language models (LLMs) often seamlessly adapt to new tasks through in-context learning (ICL) or supervised fine-tuning (SFT). However, both of these approaches face key limitations: ICL is inefficient when handling many demonstrations, and SFT incurs training overhead while sacrificing flexibility. Mapping instructions or demonstrations from context directly into adapter parameters offers an appealing alternative. While prior work explored generating adapters based on a single input context, it has overlooked the need to integrate multiple chunks of information. To address this gap, we introduce CompAs, a meta-learning framework that translates context into adapter parameters with a compositional structure. Adapters generated this way can be merged algebraically, enabling instructions, demonstrations, or retrieved passages to be seamlessly combined without reprocessing long prompts. Critically, this approach yields three benefits: lower inference cost, robustness to long-context instability, and establishes a principled solution when input exceeds the model’s context window. Furthermore, CompAs encodes information into adapter parameters in a reversible manner, enabling recovery of input context through a decoder, facilitating safety and security. Empirical results on diverse multiple-choice and extractive question answering tasks show that CompAs outperforms ICL and prior generator-based methods, especially when scaling to more inputs. Our work establishes composable adapter generation as a practical and efficient alternative for scaling LLM deployment.
@@ 1 Introduction
Large language models (LLMs) adapt to new tasks by integrating contextual information pertaining to these tasks, such as instructions, demonstrations, or extracted evidence. This adaptability is typically realized through in-context learning (ICL) (Brown et al., 2020), instruction following (Ouyang et al., 2022), and retrieval-augmented generation (Lewis et al., 2020; Borgeaud et al., 2022), where extra tokens in the prompt act as a transient memory that steers model behavior. Another common strategy is to adapt parameters directly through supervised fine-tuning (SFT), ranging from full model updates (Devlin et al., 2019; Raffel et al., 2020) to parameter-efficient variants Pfeiffer et al. (2023). While effective, both prompting and SFT face limitations. Prompt-based methods often require long contexts, which destabilize attention and degrade performance as length increases (Liu et al., 2024), while also incurring inference costs that scale with context size. SFT, on the other hand, requires additional training overhead and lacks flexibility across tasks.
@@
Figure 1: Overview of the CompAs framework. Each context $\mathbf{c}_{i}$ is mapped into an adapter by the generator $G$, while the query $\mathbf{q}$ is processed by the student $S$. The teacher $T$ processes the concatenated input $\mathbf{c};\mathbf{q}]$, and $G$ is trained so that composed adapters in $S$ align $\mathbf{y}_{s}$ with $\mathbf{y}_{t}$.
@@
To bridge this critical gap, we introduce CompAs (Compostional Adapters), a framework that maps context into adapter parameters with compositional structure at its core. See Figure 1 for a high-level overview. In CompAs, a teacher LM process concatenated contexts, while the student LM learns to approximate this behavior using only query tokens, complemented with the sum of adapters generated for individual contexts. This is facilitated by a generator network, which maps individual contexts into compositional adapters. CompAs is driven by auxiliary compositionality and reconstruction losses, which provide diagnostics for whether parameter-space operations preserve semantics and faithfulness (Jacovi and Goldberg, 2020; Turpin et al., 2023; Lampinen et al., 2022). We evaluate CompAs on multiple-choice and extractive QA tasks, where either demonstrations (MMLU, ARC-Challenge) or retrieved passages (HotpotQA) serve as context. We observe consistent improvements over ICL, fine-tuning and existing context-to-parameter methods. These gains are most pronounced when scaling to larger numbers of demonstrations, highlighting the capacity of CompAs for efficient integration of external evidence by composition in parameter-space.
@@ 2.1 Setup
Let $V$ be the vocabulary and $|V|$ its size. A language model with frozen parameters $\bm{\theta}\in\mathbb{R}^{n}$ is a function $\mathbf{f}_{\bm{\theta}}:\mathcal{X}\to\mathbb{R}^{|V|},$ where $\mathcal{X}=\Sigma^{*}$ is the set of all token sequences over the alphabet $\Sigma$. For $\mathbf{x}\in\mathcal{X}$, the output $\mathbf{f}_{\bm{\theta}}(x)\in\mathbb{R}^{|V|}$ denotes the logit vector over the vocabulary. A support context $\mathbf{c}\in\mathcal{C}$ is any textual information (e.g., a demonstration, instruction, or retrieved passage) provided alongside a query $\mathbf{q}\in\mathcal{X}$, where $\mathcal{C}\subseteq\mathcal{X}$ denotes the set of admissible contexts. We use $\mathbf{c};\mathbf{q}]\in\mathcal{X}$ to denote the concatenation of a context $\mathbf{c}$ with the query $\mathbf{q}$. Let $\Phi\subseteq\mathbb{R}^{m}$ denote the set of adapter parameters, and let $\bm{\phi}\in\Phi$ denote parameters for a single adapter, i.e., a structured modification of $\bm{\theta}$. We write $\bm{\theta}\oplus\bm{\phi}$ for the parameters obtained by composing $\bm{\theta}$ with $\bm{\phi}$.
The teacher model corresponds to the base LM applied to the concatenated input: $\mathbf{f}_{T}(\mathbf{c},\mathbf{q})\coloneqq\mathbf{f}_{\bm{\theta}}([\mathbf{c};\mathbf{q}]).$ The student model processes only the query, under parameters modified by the adapter: $\mathbf{f}_{S}^{\bm{\phi}}(\mathbf{q})\coloneqq\mathbf{f}_{\bm{\theta}\oplus\bm{\phi}}(\mathbf{q}).$ For a sequence contexts $\mathbf{c}_{1},\ldots,\mathbf{c}_{k}$, the teacher takes the concatenation $\mathbf{c}_{1};\ldots;\mathbf{c}_{k};\mathbf{q}]$, while the student uses the query together with the summed adapter $\sum_{i=1}^{k}\bm{\phi}_{i}$. For example, with two contexts, $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$, the student reduces to $\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q})=\mathbf{f}_{\bm{\theta}\oplus(\bm{\phi}_{1}+\bm{\phi}_{2})}(\mathbf{q}).$ Finally, a generator $G:\mathcal{C}\to\Phi$ maps each context $\mathbf{c}\in\mathcal{C}$ into an adapter $\bm{\phi}=G(\mathbf{c})$. In particular, for a pair of contexts $\mathbf{c_{1}}$ and $\mathbf{c_{2}}$ we define $\bm{\phi}_{1}\coloneqq G(\mathbf{c}_{1})$, $\bm{\phi}_{2}\coloneqq G(\mathbf{c}_{2})$, and $\bm{\phi}_{12}\coloneqq G([\mathbf{c}_{1};\mathbf{c}_{2}])$.
@@ 2.2 Compositionality Bound
The set of finite contexts $\mathcal{C}$, equipped with concatenation and the empty string identity ($\lambda$), forms a free monoid $(\mathcal{C},[\cdot;\cdot],\lambda)$. On the other hand, the adapter space $(\Phi,+,\mathbf{0})$ is a commutative additive monoid. A fundamental requirement for adapter compositionality is that the generator $G:\mathcal{C}\to\Phi$ be a monoid homomorphism, $G([\mathbf{c}_{1};\mathbf{c}_{2}])=G(\mathbf{c}_{1})+G(\mathbf{c}_{2})$ and $G(\lambda)=\mathbf{0},$ so that addition in the adapter space exactly mirrors concatenation in the input space. Demanding $G$ to be a homomorphism aligns these two structures. Concatenation on $\mathcal{C}$ is associative but not commutative, whereas addition on $\Phi$ is both. Thus, requiring $G$ to be a homomorphism implicitly collapses the order of contexts. This means that $G$ is a non-injective homomorphism: permuted concatenations map to the same adapter sum. We regard this loss of order as desirable in many cases, as the contribution of a demonstration or retrieved passage should often be independent of its relative position. By treating contexts as a set rather than a sequence, we aim for adapters that focus on semantics, mitigating the order sensitivity and instability observed in long-context transformers. [^1] Certain tasks may still require positional information (e.g., reasoning with ordered evidence). We show in Appendix C.3 that our framework can incorporate explicit positional encodings when needed.
@@
For a context $\mathbf{c}$ with the corresponding adapter $\bm{\phi}=G(\mathbf{c})$, the student’s discrepancy from the teacher is
$\epsilon(\mathbf{c}):=\|\mathbf{f}_{S}^{\bm{\phi}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c};\mathbf{q}])\|_{2},$ (1)
@@
$L:=\sup_{\bm{\phi}_{1},\bm{\phi}_{2},\mathbf{q}}\frac{\|\mathbf{f}_{S}^{\bm{\phi}_{1}}(\mathbf{q})-\mathbf{f}_{S}^{\bm{\phi}_{2}}(\mathbf{q})\|_{2}}{\|\bm{\phi}_{1}-\bm{\phi}_{2}\|_{2}}.$ (3)
@@
Theorem 1 (Compositionality Bound).
For any contexts $\mathbf{c}_{1},\mathbf{c}_{2}$, and query $\mathbf{q}$, with $\bm{\phi}_{i}=G(\mathbf{c}_{i})$,
$\|\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}\;\leq\;L\eta\;+\;\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}]).$
See Appendix A for the proof. The bound decomposes the student–teacher error into two interpretable sources: generator additivity error ($\eta$) and misfit on the concatenated context ($\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}])$). The Lipschitz constant $L$ propagates generator errors into the student. Thus, perfect compositionality is unattainable in practice, but approximate compositionality is achievable whenever these quantities remain small.
@@ 3.1 Context-to-Adapter Generator
As introduced in Section 2, the generator $G$ maps a context $\mathbf{c}$ to adapter parameters $\bm{\phi}=G(\mathbf{c})$. It is implemented by augmenting the base LM with additional trainable components: (i) its own LoRA adapter inserted into the LM, and (ii) a linear bottleneck followed by two projection trunks. The same frozen base LM is shared across the teacher, the student, and the generator, where only the generator’s adapter and projection components are updated during training.
@@ 3.2 Loss Function Components
Our loss function design follows Theorem 1: we reduce the student–teacher error in both single and concatenated contexts, and penalize generator non-additivity. We also include a reconstruction term that encourages adapters to faithfully encode their contexts (up to permutation), preventing collapse into trivial solutions. This auxiliary objective also regularizes training and provides a stronger learning signal, aiding optimization.
@@
We draw unlabeled queries $\mathbf{q}$ and contexts of length $k\in\{1,2\}$, letting the teacher provide soft pseudo-labels (logits). The student is trained to match these through weak supervision:
$\mathcal{L}_{\mathrm{ST}}=\;\mathbb{E}_{((\mathbf{c}_{1},\ldots,\mathbf{c}_{k}),\,\mathbf{q})}\!\left[\mathrm{KL}\!\left(\mathrm{softmax}\!\big(\mathbf{f}_{T}([\mathbf{c}_{1};\ldots;\mathbf{c}_{k};\mathbf{q}])\big)\;\big\|\;\mathrm{softmax}\!\big(\mathbf{f}_{S}^{\sum_{i=1}^{k}\bm{\phi}_{i}}(\mathbf{q})\big)\right)\right].$
where $\mathrm{KL}$ denotes Kullback–Leibler divergence. For $k=1$ this enforces single-context fidelity, while for $k=2$ it enforces compositionality.
@@
We reduce the generator additivity error $\eta$ by penalizing discrepancies between parameters generated for concatenated contexts and the sum of parameters from the individual contexts. Concretely, we draw pairs of contexts $(\mathbf{c}_{1},\mathbf{c}_{2})$ and compare generated parameters:
$\mathcal{L}_{\mathrm{ADD}}=\;\mathbb{E}_{(\mathbf{c}_{1},\mathbf{c}_{2})}\!\left[\sum_{m\in\mathcal{M}}\frac{\big\|\mathbf{W}_{m}([\mathbf{c}_{1};\mathbf{c}_{2}])-\big(\mathbf{W}_{m}(\mathbf{c}_{1})+\mathbf{W}_{m}(\mathbf{c}_{2})\big)\big\|_{F}^{2}}{\big\|\mathbf{W}_{m}([\mathbf{c}_{1};\mathbf{c}_{2}])\big\|_{F}^{2}+\delta}\right],$
where $\mathbf{W}_{m}(\mathbf{c})\in\{\mathbf{A}_{m}(\mathbf{c}),\mathbf{B}_{m}(\mathbf{c})\}$ denotes either of the two LoRA matrices generated for module $m\in\mathcal{M}$, $\|\cdot\|_{F}$ is the Frobenius norm, and $\delta>0$ is a small constant for stability.
@@
To encourage faithfulness and improve training stability, we require adapters to recoverably encode the information contained in their contexts. We introduce a special query token [RECON], which prompts the student to reconstruct the support context from its adapter in an autoregressive manner. This discourages collapse to trivial solutions and provides an auxiliary learning signal through the cross-entropy (CE) loss:
$\mathcal{L}_{\mathrm{RECON}}=\;\mathbb{E}_{\mathbf{c}}\!\left[\mathrm{CE}\!\big(\mathbf{c},\,\mathbf{f}_{S}^{\bm{\phi}}(\texttt{[RECON]})\big)\right].$
@@
Hyperparameter choices and training details are provided in Appendix B.

### v3

@@ Abstract
Large language models (LLMs) often seamlessly adapt to new tasks through in-context learning (ICL) or supervised fine-tuning (SFT). However, ICL is inefficient when handling many demonstrations, and SFT incurs training overhead while sacrificing flexibility. Mapping instructions or demonstrations from context directly into adapter parameters offers an appealing alternative. While prior work explored generating adapters based on a single input context, it has overlooked the need to integrate multiple chunks of information. To address this gap, we introduce CompAs, a meta-learning framework that translates context into adapter parameters with a compositional structure that allows them to be merged algebraically. This approach yields three benefits: lower inference cost, improved stability under long contexts, and establishes a principled solution when input exceeds the model’s context window. Furthermore, CompAs reversibly encodes information into adapter parameters, enabling recovery of the original input context and facilitating safety. Empirical results on diverse multiple-choice and extractive question answering tasks show that CompAs outperforms ICL and prior generator-based methods, especially when scaling to more inputs. Our work establishes composable adapter generation as a practical and efficient alternative for scaling LLM deployment.
LLMs
@@ 1 Introduction
Large language models (LLMs) adapt to new tasks by integrating contextual information pertaining to these tasks, such as instructions, demonstrations, or extracted evidence. This adaptability is typically realized through in-context learning (ICL) (Brown et al., 2020), instruction following (Ouyang et al., 2022), and retrieval-augmented generation (Lewis et al., 2020; Borgeaud et al., 2022), where extra tokens in the prompt act as a transient memory that steers model behavior. Another common strategy is to adapt parameters directly through supervised fine-tuning (SFT), ranging from full model updates (Devlin et al., 2019; Raffel et al., 2020) to parameter-efficient variants (Pfeiffer et al., 2023). While effective, both prompting and SFT face limitations. Prompt-based methods often require long contexts, which destabilize attention and degrade performance as length increases (Liu et al., 2024), while also incurring inference costs that scale with context size. SFT, on the other hand, requires additional training overhead and lacks flexibility, as an individual model is trained for each task.
@@
Figure 1: Overview of the CompAs framework. Each context $\mathbf{c}_{i}$ is mapped into an adapter by the generator $G$, while the query $\mathbf{q}$ is processed by the student $S$. The teacher $T$ processes the concatenated input $\mathbf{c};\mathbf{q}]$, and $G$ is trained so that composed adapters in $S$ align $\mathbf{y}_{s}$ with $\mathbf{y}_{t}$, effectively transferring contextual information from input space to parameter space.
@@
To bridge this critical gap, we introduce CompAs (Compositional Adapters), a framework that maps context into adapter parameters with compositional structure at its core. See Figure 1 for a high-level overview. In CompAs, a teacher LM process concatenated contexts, while the student LM learns to approximate this behavior using only query tokens, complemented with the sum of adapters generated for individual contexts. This is facilitated by a generator network, which maps individual contexts into compositional adapters. CompAs is driven by auxiliary compositionality and reconstruction losses, which provide diagnostics for whether parameter-space operations preserve semantics and faithfulness (Jacovi and Goldberg, 2020; Turpin et al., 2023; Lampinen et al., 2022). We evaluate CompAs on multiple-choice and extractive QA tasks, where either demonstrations (MMLU, ARC-Challenge) or retrieved passages (HotpotQA) serve as context. We observe consistent improvements over ICL, fine-tuning and existing context-to-parameter methods. These gains are most pronounced when scaling to larger numbers of demonstrations, highlighting the capacity of CompAs for efficient integration of external evidence by composition in parameter-space.
@@ 2.1 Setup
Let $V$ be the vocabulary and $|V|$ its size. A language model with frozen parameters $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is a function $\mathbf{f}_{\boldsymbol{\theta}}:\mathcal{X}\to\mathbb{R}^{|V|},$ where $\mathcal{X}=\Sigma^{*}$ is the set of all token sequences over the alphabet $\Sigma$. For $\mathbf{x}\in\mathcal{X}$, the output $\mathbf{f}_{\boldsymbol{\theta}}(x)\in\mathbb{R}^{|V|}$ denotes the logit vector over the vocabulary. A support context $\mathbf{c}\in\mathcal{C}$ is any textual information (e.g., a demonstration, instruction, or retrieved passage) provided alongside a query $\mathbf{q}\in\mathcal{X}$, where $\mathcal{C}\subseteq\mathcal{X}$ denotes the set of admissible contexts. We use $\mathbf{c};\mathbf{q}]\in\mathcal{X}$ to denote the concatenation of a context $\mathbf{c}$ with the query $\mathbf{q}$.
Let $\Phi\subseteq\mathbb{R}^{m}$ denote the set of adapter parameters, and let $\boldsymbol{\phi}\in\Phi$ denote parameters for a single adapter, i.e., a structured modification of $\boldsymbol{\theta}$. We write $\boldsymbol{\theta}\oplus\boldsymbol{\phi}$ for the parameters obtained by composing $\boldsymbol{\theta}$ with $\boldsymbol{\phi}$.
We now formalize the teacher–student setup that underlies our framework. The teacher model corresponds to the base language model applied to the concatenated input:
$\mathbf{f}_{T}(\mathbf{c},\mathbf{q})\coloneqq\mathbf{f}_{\boldsymbol{\theta}}\!\left([\mathbf{c};\mathbf{q}]\right).$
The student model processes only the query, using parameters modified by an adapter:
$\mathbf{f}_{S}^{\boldsymbol{\phi}}(\mathbf{q})\coloneqq\mathbf{f}_{\boldsymbol{\theta}\oplus\boldsymbol{\phi}}(\mathbf{q}).$
Throughout this work, both teacher and student share the same frozen base parameters $\boldsymbol{\theta}$, ensuring that any behavioral differences arise solely from the presence or absence of adapter-based context encoding. While the formulation could accommodate distinct base models in principle, we deliberately keep the underlying model fixed to isolate the effect of adapters.
For a sequence of contexts $\mathbf{c}_{1},\ldots,\mathbf{c}_{k}$ (denoted compactly by $\mathbf{c}_{1:k}$), the teacher receives the concatenated input $\mathbf{c}_{1:k};\mathbf{q}]$, while the student processes only $\mathbf{q}$ together with the summed adapter parameters $\sum_{i=1}^{k}\boldsymbol{\phi}_{i}$. For example, with two contexts $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$, the student model becomes
$\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q})=\mathbf{f}_{\boldsymbol{\theta}\oplus(\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2})}(\mathbf{q}).$
Finally, a generator $G:\mathcal{C}\to\Phi$ maps each context $\mathbf{c}\in\mathcal{C}$ to an adapter $\boldsymbol{\phi}=G(\mathbf{c})$. In particular, for two contexts $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$, we define
$\boldsymbol{\phi}_{1}=G(\mathbf{c}_{1}),\quad\boldsymbol{\phi}_{2}=G(\mathbf{c}_{2}),\quad\boldsymbol{\phi}_{12}=G([\mathbf{c}_{1};\mathbf{c}_{2}]).$
@@ 2.2 Compositionality Bound
The set of finite contexts $\mathcal{C}$, equipped with concatenation and the empty string identity ($\lambda$), forms a free monoid $(\mathcal{C},[\cdot;\cdot],\lambda)$. On the other hand, the adapter space $(\Phi,+,\mathbf{0})$ is a commutative additive monoid. A fundamental requirement for adapter compositionality is that the generator $G:\mathcal{C}\to\Phi$ is a monoid homomorphism, $G([\mathbf{c}_{1};\mathbf{c}_{2}])=G(\mathbf{c}_{1})+G(\mathbf{c}_{2})$ and $G(\lambda)=\mathbf{0}.$ This ensures that addition in the adapter space exactly mirrors concatenation in the input space. Demanding $G$ to be a homomorphism aligns these two structures. Concatenation on $\mathcal{C}$ is associative but not commutative, whereas addition on $\Phi$ is both. Thus, requiring $G$ to be a homomorphism implicitly collapses the order of contexts. This means that $G$ is a non-injective homomorphism: permuted concatenations map to the same adapter sum. We regard this loss of order as desirable in many cases, as the contribution of a demonstration or retrieved passage should often be independent of its relative position. By treating contexts as a set rather than a sequence, we aim for adapters that focus on semantics, mitigating the order sensitivity and instability observed in long-context transformers. [^1] Certain tasks may still require positional information (e.g., reasoning with ordered evidence). We show in Appendix C.3 that our framework can incorporate explicit positional encodings when needed.
@@
For a context $\mathbf{c}$ with the corresponding adapter $\boldsymbol{\phi}=G(\mathbf{c})$, the student’s discrepancy from the teacher is
$\epsilon(\mathbf{c}):=\|\mathbf{f}_{S}^{\boldsymbol{\phi}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c};\mathbf{q}])\|_{2},$ (1)
@@
$L:=\sup_{\boldsymbol{\phi}_{1},\boldsymbol{\phi}_{2},\mathbf{q}}\frac{\|\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}}(\mathbf{q})-\mathbf{f}_{S}^{\boldsymbol{\phi}_{2}}(\mathbf{q})\|_{2}}{\|\boldsymbol{\phi}_{1}-\boldsymbol{\phi}_{2}\|_{2}}.$ (3)
@@
Theorem 2.1 (Compositionality Bound).
Let $\boldsymbol{\phi}=G(\mathbf{c})$ denote the adapter parameters generated from a context $\mathbf{c}$. Let $\epsilon$ denote the teacher–student compositional error, $\eta$ the generator additivity error, and $L$ the sensitivity of the student with respect to adapter parameters. Then, for any contexts $\mathbf{c}_{1},\mathbf{c}_{2}$ and query $\mathbf{q}$,
$\|\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}\;\leq\;L\eta\;+\;\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}]).$
See Appendix A for the proof. The bound decomposes the student–teacher error into two interpretable sources: generator additivity error $\eta$ and misfit on the concatenated context $\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}])$. The Lipschitz constant $L$ propagates generator errors into the student. Thus, perfect compositionality is unattainable in practice, but approximate compositionality is achievable whenever these quantities remain small.
@@ 3.1 Context-to-Adapter Generator
As introduced in Section 2, the generator $G$ maps a context $\mathbf{c}$ to adapter parameters $\boldsymbol{\phi}=G(\mathbf{c})$. It is implemented by augmenting the base LM with additional trainable components: (i) its own LoRA adapter inserted into the LM, and (ii) a linear bottleneck followed by two projection trunks. The same frozen base LM is shared across the teacher, the student, and the generator, where only the generator’s adapter and projection components are updated during training.
@@ 3.2 Loss Function Components
Our loss function design follows Theorem 2.1: we reduce the student–teacher error in both single and concatenated contexts, and penalize generator non-additivity. We also include a reconstruction term that encourages adapters to faithfully encode their contexts (up to permutation), preventing collapse into trivial solutions. This auxiliary objective also regularizes training and provides a stronger learning signal, aiding optimization.
@@
We draw unlabeled queries $\mathbf{q}$ together with $k$ contexts, and obtain soft pseudo-labels from the teacher. For brevity, we introduce the shorthands $\mathbf{p}_{T}(\mathbf{c}_{1:k},\mathbf{q})$ and $\mathbf{p}_{S}^{\boldsymbol{\phi}}(\mathbf{q})$, denoting the teacher and student predictive distributions, respectively. Concretely, we define
$\mathbf{p}_{T}(\mathbf{c}_{1:k},\mathbf{q})$ $=\mathrm{softmax}\!\big(\mathbf{f}_{T}([\mathbf{c}_{1:k};\mathbf{q}])\big),$
$\mathbf{p}_{S}^{\boldsymbol{\phi}}(\mathbf{q})$ $=\mathrm{softmax}\!\big(\mathbf{f}_{S}^{\sum_{i=1}^{k}\boldsymbol{\phi}_{i}}(\mathbf{q})\big).$
The student is trained to match the teacher’s predictive distribution via weak supervision by minimizing the Kullback–Leibler divergence, denoted by $D_{\mathrm{KL}}(\cdot\,\|\,\cdot)$, between their output distributions:
$\mathcal{L}_{\mathrm{ST}}=\mathbb{E}_{(\mathbf{c}_{1},\ldots,\mathbf{c}_{k}),\,\mathbf{q}}\bigg[D_{\mathrm{KL}}\!\Big(\mathbf{p}_{T}(\mathbf{c}_{1:k},\mathbf{q})\;\big\|\;\mathbf{p}_{S}^{\boldsymbol{\phi}}(\mathbf{q})\Big)\bigg].$
Our overall objective considers $k\in\{1,2\}$, where $k=1$ enforces fidelity to a single context, while $k=2$ encourages compositional consistency across two contexts.
@@
We reduce the generator additivity error $\eta$ by penalizing discrepancies between parameters generated for concatenated contexts and the sum of parameters from the individual contexts. Concretely, we draw pairs of contexts $(\mathbf{c}_{1},\mathbf{c}_{2})$ and compare generated parameters. For each module $m\in\mathcal{M}$, let $\mathbf{W}_{m}(\mathbf{c})\in\{\mathbf{A}_{m}(\mathbf{c}),\mathbf{B}_{m}(\mathbf{c})\}$ denote one of the two LoRA matrices generated from context $\mathbf{c}$. We write $\|\cdot\|_{F}$ for the Frobenius norm and let $\delta>0$ be a small stability constant. We define the normalized additivity discrepancy
$\Delta_{m}(\mathbf{c}_{1},\mathbf{c}_{2})=\frac{\left\|\mathbf{W}_{m}([\mathbf{c}_{1:2}])-\mathbf{W}_{m}(\mathbf{c}_{1})-\mathbf{W}_{m}(\mathbf{c}_{2})\right\|_{F}^{2}}{\left\|\mathbf{W}_{m}([\mathbf{c}_{1:2}])\right\|_{F}^{2}+\delta},$
and minimize the additivity objective
$\mathcal{L}_{\mathrm{ADD}}=\mathbb{E}_{(\mathbf{c}_{1},\mathbf{c}_{2})}\left[\sum_{m\in\mathcal{M}}\Delta_{m}(\mathbf{c}_{1},\mathbf{c}_{2})\right].$
@@
To encourage faithfulness and improve training stability, we require adapters to recoverably encode the information contained in their contexts. To this end, we introduce a special query token [RECON], which prompts the student to reconstruct the support context $\mathbf{c}=(c_{1},\ldots,c_{T})$ autoregressively from its adapter state, where $c_{t}$ denotes the token at position $t$ and $\mathbf{c}_{<t}=(c_{1},\ldots,c_{t-1})$ the prefix preceding position $t$. Let $p_{S}^{\boldsymbol{\phi}}(v\mid\mathbf{s})$ denote the student’s conditional predictive distribution over the next token $v$ given an arbitrary conditioning sequence $\mathbf{s}$. Under the reconstruction prompt, the conditioning sequence is $\mathbf{s}=\left[\texttt{[RECON]};\mathbf{c}_{<t}\right]$. This discourages collapse to trivial solutions and provides an auxiliary learning signal through the cross-entropy loss:
$\mathcal{L}_{\mathrm{RECON}}=-\mathbb{E}_{\mathbf{c}}\left[\sum_{t=1}^{T}\log p_{S}^{\boldsymbol{\phi}}(c_{t}\mid\texttt{[RECON]};\mathbf{c}_{<t})\right].$
@@
The coefficients $\lambda_{\text{ST}},\lambda_{\text{ADD}},$ and $\lambda_{\text{RECON}}$ balance alignment, additivity regularization, and reconstruction fidelity, respectively. Together, these terms support compositional generalization while maintaining faithful encoding of contextual information. Hyperparameter choices and training details are provided in Appendix B.